OpenStudy (automaticloveletter):

What would most likely be the approximate height of the plant after 7 weeks? 4.8 centimeters 14.4 centimeters 16.5 centimeters 19.5 centimeters

OpenStudy (automaticloveletter):

http://prntscr.com/7uzkkn

OpenStudy (triciaal):

one approach find the slope of the line change in y/ change in x slope is the same for any 2 points on that line using one of the points and (7, y) solve for y the height when x = 7

OpenStudy (automaticloveletter):

@triciaal I'm confused

OpenStudy (automaticloveletter):

@ChillOut

OpenStudy (chillout):

We have a linear equation. The form of a linear equation is given by y= mx + b, where "m" is the angular coefficient and "b" the y intercept. "b" is given by \(\frac{ \Delta y }{ \Delta x }\)

OpenStudy (chillout):

Moving on... "b", the "y intecept" is 2 as we can see from graph 

OpenStudy (chillout):

So... now to find the coefficient "m". We need to use known points, and the easiest ones for that are (0,2) and (5,4). We do \(\frac{ \Delta y }{ \Delta x }\) and we have \[\frac{ 4-2 }{ 5-0 }=\frac{ 2 }{ 5 }\] Now we need to plug m and b in the equation... \[\frac{ 2 }{ 5 }x +2\] This is the equation for this line. After this, we plug 7 in the "x".

OpenStudy (chillout):

And we will have \[\frac{ 2 }{ 5 }*7 +2=\frac{14}{5}+2 = \frac{24}{5}=4.8 cm\]

OpenStudy (chillout):

Did you get it? :)

OpenStudy (automaticloveletter):

Got it

OpenStudy (chillout):

All right then. If you need any help I'll be back later. Just ask ;)

OpenStudy (chillout):

Oh, and in my first response it should be "m" is given by dy/dx. major mistake there by me.

OpenStudy (triciaal):

@automaticloveletter you are confused on finding the slope of the line even after telling you it's the change in y / change in x ?

OpenStudy (automaticloveletter):

@triciaal Yeah, well I got it now. Thanks anyways.